OpenStudy (twistrose):

How can I find the answer to the question? 1 ---- 3^-4 Please help. I don't understand.

OpenStudy (sissyedgar):

Well.... For a problem like.. \[3^{-4}\] My teacher always taught me to use the calculator so I don't know how to do it by hand. But I do know how to find the answer on a scientific calculator. Srry.

OpenStudy (twistrose):

its ok and the one is included

jhonyy9 (jhonyy9):

use x^-1 = 1/x

jhonyy9 (jhonyy9):

so 3^-4 = 1/3^4 but there are 1 1 ------ = ----- = ? 3^-4 1 ----- 3^4

jhonyy9 (jhonyy9):

hope you know that a -- b ---- = a/b * d/c yes ? c -- d

OpenStudy (mhchen):

\[\frac{ 1 }{ 3^{-4} } = 3^{4} = 81\]

jhonyy9 (jhonyy9):

dont give directly right answers bc, this is opposite to OS rules @TheSmartOne